### v5

# Designing quantum chemistry algorithms with just-in-time compilation

Abstract
We introduce just-in-time (JIT) compilation to the integral kernels for Gaussian-type orbitals (GTOs) to enhance the efficiency of electron repulsion integral computations. For Coulomb and exchange (JK) matrices, JIT-based algorithms yield a 2× speedup for the small 6-31G* basis set over GPU4PySCF v1.4 on an NVIDIA A100-80G GPU. By incorporating a novel algorithm designed for orbitals with high angular momentum, the efficiency of JK evaluations with the large def2-TZVPP basis set is improved by up to 4×. The core CUDA implementation is compact, comprising only ~1,000 lines of code, including support for single-precision arithmetic. Furthermore, the single-precision implementation achieves a 3× speedup over the previous state-of-the-art.
[…]
1 Introduction
Traditional quantum chemistry software [16, 33, 30] has historically relied on ahead-of-time (AOT) compilation and monolithic code generation, with tightly coupled Fortran or C++ implementations [21, 8, 10] tuned for a fixed set of use cases. Although this design has served well for decades, it poses serious limitations in some modern high-performance computing (HPC) environments where GPUs are the primary source of computational power. Algorithms in electronic structure theory often require handling a combinatorial variety of tensor contraction patterns and sparsity conditions that depend on the input molecular system. Static compilation must conservatively support all conditional branches, hyperparameter combinations, and numerical thresholds, leading to bloated binaries, excessive control flow divergence, poor cache and register utilization, and inefficient use of memory bandwidth. On GPUs, these inefficiencies are further magnified by strict constraints on register pressure, warp divergence, and hardware occupancy, often requiring significant manual optimization. For example, most quantum chemistry packages deal with low-angular-momentum integrals successfully but still struggle with the performance of high-angular momentum integrals, which suffer from all of these problems.
Just-in-time (JIT) compilation offers a transformative solution by enabling runtime code specialization based on the actual inputs: e.g., integral angular-momentum pattern, the number of primitives in an orbital shell, and task type. By generating code on demand, JIT techniques can eliminate irrelevant branches, tailor memory access patterns, and fuse computational stages that would be separate in statically compiled workflows. In GPU architectures, this dynamic specialization translates into reduced register pressure, improved memory coalescing, and better occupancy. Similar ideas have seen widespread adoption in the machine learning community, where frameworks such as PyTorch’s TorchScript [24], Triton [28] and JAX [5] leverage JIT to fuse operations, reduce memory movement, and generate hardware-specific kernels. These successes demonstrate the maturity and impact of JIT compilation as a general-purpose performance strategy in high-throughput numerical computing, and motivate its application to quantum chemistry workloads.
Beyond runtime performance, just-in-time compilation also offers substantial advantages to developers. JIT facilitates rapid prototyping and testing: since code is generated at runtime, one can easily benchmark different kernel variants, tiling strategies, or instruction sequences, without recompiling and linking the entire application. This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures. Furthermore, the developers are able to design novel algorithms which are not general enough under the AOT framework. In this paper, we demonstrate several algorithms due to this benefit. The development experience can be even further improved: Within a narrow area of application such as machine learning, SQL and GPU programming, one can design domain-specific languages (DSLs) via LLVM [15]. Frameworks like TensorFlow[1], PyTorch[24], JAX[5], and TVM incorporate their own DSLs—such as TensorFlow’s GraphDef and XLA HLO, PyTorch’s TorchScript and FX, JAX’s JAXpr, and TVM’s TIR—to represent algorithms in intermediate forms that enable aggressive optimization and efficient execution. These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
JIT compilation benefits mixed precision algorithms in quantum chemistry, which naturally accelerate electron repulsion integrals by exploiting the Schwarz inequality [17, 20]. Most tasks can be calculated in single-precision (FP32) as compared to the commonly-used double-precision (FP64). Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads. These benefits can be better utilized by JIT compilation without significant changes to the code.
Contributions: 1) We introduce JIT to the quantum chemistry community and optimize the two-electron integral kernels (1q1t in Sec. 3.2) with compilation techniques. Our framework integrates seamlessly with the quantum chemistry package GPU4PySCF [16, 33] and is released as the open-source JoltQC library. 2) We develop a novel fragmentation algorithm for high angular momentum integrals, improving data locality and alleviating memory-bandwidth bottlenecks through multilevel reduction. 3) We benchmark single-precision integral kernels under our JIT framework and demonstrate substantial performance gains when combining JIT compilation with our new algorithmic strategies.
2 A short formulation of Rys quadrature
Evaluation of the electron repulsion integral is one major bottleneck of density functional theory (DFT), the most widely used electronic structure theory [4]. In this paper, the two-electron repulsion integrals $(ij|kl)$ are calculated with the Rys quadrature algorithm [9]. Other popular integral algorithms such as McMurchie–Davidson [19] and Head-Gordon-Pople [11] have been implemented in [2, 30, 29]. Reimplementation of these methods with JIT is out of the scope of this paper. In this section, we briefly formulate the Rys quadrature algorithm for clarifying the notations and data dependency used in the rest of the paper. For detailed derivation and numerical stability analysis, we refer the readers to [9].
Here we are using ijklijkl notations instead of conventional $\mu\nu\lambda\sigma$ notations to index atomic orbitals, to be consistent with the pseudo-code later in this paper. Since each of the Gaussian type orbitals is contracted from a set of primitive Gaussian functions [12], the integral tensor for a shell quartet is expressed as the sum of the primitive integral tensors
$(ij|kl)=\sum_{pqrs}c_{p}c_{q}c_{r}c_{s}(pq|rs)$ (1)
$c_{p}$, $c_{q}$}, $c_{r}$, and $c_{s}$ are the pre-defined contraction coefficients for each shell in the basis set. The number of Gaussian functions in each primitive shell is $l(l+1)/2$ or $2l+1$ for cartesian atomic orbitals or spherical atomic orbitals respectively, where $l$ is the angular momentum of the shell. Only Cartesian orbitals are considered in the following sections, and spherical orbitals are just a linear combination of Cartesian orbitals with fixed transformation coefficients. These linear transformations can be performed efficiently outside the kernel. As a result, each dimension of the primitive integral $(pq|rs)$ is $N^{f}=l(l+1)/2$. We denote the dimension of the integral for each shell quartet as $(N_{i}^{f},N_{j}^{f},N_{k}^{f},N_{l}^{f})$.
Within Rys quadrature formalism, each primitive integral can be evaluated as a weighted sum of the integrant at specific quadrature root points
$(pq|rs)=\sum_{r}^{N_{r}}(pq|rs)(t_{r})\cdot w_{r}\\$ (2)
where $N_{r}$ is the number of Rys roots. $t_{r}$ and $w_{r}$ are the roots and weights of Rys quadrature. The roots and weights depend on the exponents and center (atomic) positions of the four primitive Gaussian functions, the total angular momentum, and a range-separation parameter $\omega$ if needed. The calculation of Rys roots and weights relies on interpolations from tabulated data, and the details have been documented in [9]. The number of Rys roots is determined at the compilation time, and only the data[33, 26] associated with the number of roots are compiled into the binary. The primitive integral at each root position can be factored into intermediate variables in three spatial directions:
$(pq|rs)(t_{r})=$ $I_{x}(a_{x},b_{x},c_{x},d_{x},t_{r})\cdot I_{y}(a_{y},b_{y},c_{y},d_{y},t_{r})$ (3)
$\cdot I_{z}(a_{z},b_{z},c_{z},d_{z},t_{r})$
for any $(p,a_{x},a_{y},a_{z})\in S_{l_{i}},(q,b_{x},b_{y},b_{z})\in S_{l_{j}},$
$(r,c_{x},c_{y},c_{z})\in S_{l_{k}},(s,d_{x},d_{y},d_{z})\in S_{l_{l}}$
$I_{x}$, $I_{y}$, and $I_{z}$ are the intermediate variables introduced in [7, 9]. Each of them is a tensor of shape $(l_{i}+1,l_{j}+1,l_{k}+1,l_{l}+1)$. They depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only. They can be efficiently evaluated using recursive relations, documented in [9]. But storing those values on the limited resources on GPU can be tricky, and several strategies will be discussed in Sec. 3.1 and Sec. 3.2. The index set for enumerating the Guassian functions of a primitive shell is defined as
$S_{l}:=\{(i,a,b,c)|a+b+c=l,i=\texttt{idx}(a,b,c)\}$
The idx function is defined differently in various quantum chemistry packages. One will need to double check the conventions to align the quantities in atomic basis, such as density matrix, molecular orbitals, and so on. We adopt the convention of reversed lexicographic order. The indexing operation in general makes the memory access random and less efficient.
Because of its huge size, the four-dimensional integral tensor is not formed explicitly in the main storage, but instead contracted with the density matrix $D$ to form the Coulomb ($J$) and exchange ($K$) potential matrices:
$J_{ij}=\sum_{kl}(ij|kl)D_{kl},\text{ and }K_{ik}=\sum_{jl}(ij|kl)D_{jl}$ (4)
For a molecular system with real-valued orbitals, the 8-fold symmetry of $(ij|kl)$ can be used to save computational cost. In particular, for the integral element $(ij|kl)$ composed of the $i$-th, $j$-th, $k$-th and $l$-th atomic orbital, it is contracted with two density matrix elements $D_{kl}$ and $D_{lk}$ to form two matrix elements $J_{ij}$ and $J_{ji}$, and it is contracted with four density matrix elements $D_{jl}$, $D_{jk}$. $D_{il}$, $D_{ik}$ to form four $K$ matrix elements $K_{ik}$, $K_{il}$, $K_{jk}$, $K_{jl}$.
Conventionally, the Schwarz inequality for shell quartets
$|(ij|kl)|\leq\max_{i\in S_{l_{i}},j\in S_{l_{j}}}\sqrt{(ij|ij)}\cdot\max_{k\in S_{l_{k}},l\in S_{l_{l}}}\sqrt{(kl|kl)}$ (5)
is used to screen the significant shell quartets, i.e. all integrals within an insignificant shell quartet are considered negligible and not evaluated. The cost of generating the significant shell quartet is almost negligible, even for the low-angular-momentum integrals. Other screening techniques can be combined with the algorithms. They will be discussed in section 6.
In the direct self-consistent field (SCF) algorithm, the integrals are further screened by the density matrix
$|\Delta J_{ij}|\leq\sqrt{(ij|ij)}\sum_{kl}|\Delta D_{kl}|\sqrt{(kl|kl)}$ (6)
and
$|\Delta K_{ik}|\leq\sum_{jl}|\Delta D_{jl}|\sqrt{(ij|ij)}\sqrt{(kl|kl)}$ (7)
where $\Delta D$, $\Delta J$, and $\Delta K$ are the changes of density matrix, Coulomb potential, and exchange potential between two steps of SCF iterations (i.e. incremental SCF), respectively. Since $\Delta D$ is getting smaller along the SCF iterations, the screening is more effective. In practice, the density matrix screening is performed at shell pair level, which is similar to Eq. (5).
3 Design and Algorithms
The input of a quantum chemistry calculation contains the protocol of the method, geometry of a molecule, and various thresholds. These inputs are unknown when a typical quantum chemistry package is compiled. Therefore, all types of input have to be handled as dynamic variables. The idea of JIT design is to feed the input information to the compiler at compilation time such that the compiler can aggressively optimize the code. On the other hand, if the compilation completely depends on all the input information, i.e. if we recompile for every job even if successively jobs are only different in geometry, the recompilation may slow down the overall execution time, since the compilation can be more expensive than the actual calculation with an improper design. In the design of JIT programs, distinguishing between static and dynamic variables is crucial for user experience and performance. In general, we follow this principle: The information of quantum chemistry protocol and device information are static for compilation. That is, when the quantum chemistry method, basis set, or type of GPUs is changed, the kernels should be recompiled. The same kernels will be used for different molecules. All the molecule information, including the atomic numbers, atomic types and atomic coordinates, and the total charge and spin of the system, are dynamic variables for compilation.
More specifically, we select the angular momentum and contraction pattern of the basis set, as well as the task type, as the template parameters for the kernels. They are described in Table 1. Once these variables are fixed, the size of any intermediate variable in the integral kernels is known at compilation time. If those parameters are changed, we have to recompile the kernels. With such choices, we avoid complicated runtime logic flows for different data type, task type, and unrolling strategies. Thus the amount of source code is minimal: only ~1,000 lines of code for each algorithm, much less than the ~20,000 lines in GPU4PySCF v1.4.
Template
Parameter
Data Type Description
$(l_{i},l_{j},l_{k},l_{l})$ int Angular momentum of shells
do_j bool Whether to compute the J matrix
do_k bool Whether to compute the K matrix
rys_type int 0: no range separation (full range),
1: long-range, -1: short-range
n_dm int # of density matrices
(e.g. two for unrestricted DFT)
DataType type float or double precision
$(n_{i}^{p},n_{j}^{p},n_{k}^{p},n_{l}^{p})$ int # of primitives in the
(i, j, k, l) shells
$(F_{i},F_{j},F_{k},F_{l})$ int Fragment sizes of the
$(i,j,k,l)$ shells (details in Sec. 3.2)
Table 1: Static parameters for integral kernel generation.
The parameters of the basis set are dynamic variables. Since many elements share the same contraction pattern but different coefficients and exponents, we do not need to compile multiple kernels for the same contraction pattern. Generally, reading those data is not the bottleneck of the kernels. All of the dynamic parameters are summarized in 2.
Input
Data Type Description
$(c_{i},c_{j},c_{k},c_{l})$ double/single contraction coefficients in basis
$(e_{i},e_{j},e_{k},e_{l})$ double/single exponents in basis
coords double/single coordinates of atoms
$D_{ij}$ double density matrix as input
$J_{ij}$ double coulomb potential as output
$K_{ij}$ double exchange potential as output
$N_{ao}$ int # of atomic orbitals
$\omega$ double range-separation parameter
Table 2: Dynamic parameters as the input of integral kernels. The data type of basis parameters are determined by the static variable DataType.
An example of the DFT self-consistent field (SCF) workflow is presented in Fig. 1. The workflow employs NVRTC [22], the runtime compilation library for CUDA, via CuPy[23] as JIT compiler. NVRTC offers the direct control over shared memory, as well as a relatively fast compilation. Other compilers, such as Numba, can be employed similarly with minimal modification, which will be discussed in Sec. 6
Figure 1: Workflow of SCF iterations with NVRTC as JIT compiler.
3.1 One quartet one thread (1q1t) algorithm
The first algorithm we present is largely the same as the classical Rys quadrature algorithm on CPUs [26], except for a more aggressive loop unrolling strategy is used. First, a list of significant shell quartets is generated by running over all possible combinations screened by the Schwartz inequality. Each quartet is assigned to one thread for evaluating all elements of the integral in Eq. (1) (including all primitives in a contraction and all functions in a shell), and contract it with the density matrix according to Eq. (4), until the pool of all significant shell quartet is empty. This algorithm was implemented in the early version of GPU4PySCF (v1.0 and before). This straightforward implementation has two performance issues. 1) There are still multiple loops not unrolled due to unknown variables at compilation time, such as number of primitivies for each shell and number of density matrices. This is actually the main bottleneck of low-angular-momentum integrals. 2) Significant register spill out onto local memory for high angular momentum. In this subsection, we will show that with more known variables at compilation time, the performance is significantly improved. The second issue will be discussed in the next subsection.
With the design in this work, when handling one shell quartet, the ranges of all loops are known at compilation time. The JIT compiler can aggressively unroll those loops, especially loops for primitives, and use registers more efficiently.
In the single-precision implementation, the exponents, contraction coefficients, and atomic coordinates are prepared and loaded in FP32. The density matrix values are loaded as FP64 and then casted to FP32 in kernels. Inside the kernel, all the arithmetic operations are performed in FP32 and accelerated with the --use_fast_math compiler flag. One should notice that fast-math intrinsics for single precision is much faster, but not IEEE-754 accurate. When $J$ and $K$ are calculated, they are first casted to FP64, then accumulated with atomicAdd in FP64.
We benchmark the performance of JK kernels with a straight peptide chain composed of 120 glycines (gly120), and 6-31G [6] basis set. Density matrix screening is turned off for profiling purpose. The timings on A100-80G and A10-24G are shown in Figure 2 and 3 respectively. The overall speedup of FP64 against AOH compilation in GPU4PySCF v1.4 is about 2.0x. For low-angular-momentum, the evaluations of exp and erf functions are often the computational bottleneck. Thanks to the special hardware acceleration of these functions on GPU, the overall speedup of FP32 reaches about 5-10x on A100-80G. Since A10-24G has 1/32 hardware units for FP64 compared to FP32, FP64 algorithms are always bounded by the limited compute units. The improvement with JIT is marginal. When switched to FP32, the 1q1t algorithm is accelerated by 30x.
There are still several corner-case kernels. When the number of primitives is greater than 5, the NVRTC compiler refuses to unroll the nested loops, even if we specify #pragma unroll. Performance drop is observed. The issue will be discussed in Appendix (Section 8). Another case is $(pp|pp)$ in double precision which requires 96 registers per thread for the intermediate variables, and 162 registers for the integral elements. The kernel consumes at least 258 registers, which already exceeds the hard limit per thread (255 registers) by the CUDA platform. The register spill affects the performance dramatically. Higher angular momentum integrals suffer from the same issues. They will be handled by different algorithm later in Section 3.2.
Figure 2: Averaged kernel runtime of JoltQC v0.1 JK calculation using 1q1t algorithm, on the gly120 system, 6-31G basis set. One Nvidia A100-80G GPU is used. The performance is compared with GPU4PySCF v1.4. Top 12 most time-consuming kernels are selected, and are labeled as (angular momentum combination)/(contraction pattern combination).
Figure 3: Averaged kernel runtime of JoltQC v0.1 JK calculation using 1q1t algorithm, on the gly120 system, 6-31G basis set. One Nvidia A10-24G GPU is used. The performance is compared with GPU4PySCF v1.4. Top 12 most time-consuming kernels are selected, and are labeled as (angular momentum combination)/(contraction pattern combination).
3.2 Quartet fragmentation (1qnt) algorithm
For high-angular momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers. From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy is introduced to split the integral into multiple threads. Instead of one block thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet. The strategy significantly improved the performance of the high-angular integrals. In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory). Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variable $I_{x}$, $I_{y}$ and $I_{z}$ will be stored in shared memory. Then, each thread in the group will compute and store into registers only a fraction of the integrals. For example, the shape of the integral (dd|dd) is (6,6,6,6), each thread computes a fragment with the size of (3,3,3,3). 16 threads are needed to store the entire (dd|dd) integral. In the reduction stage, each thread is responsible for a fragment of 3x3 $v_{j}$ and $v_{k}$. The fragments $v_{j}$ and $v_{k}$ are further reduced at the thread block level. The reduction scheme is presented in Algorithm 1. The shape of the intermediate variables are also known at compilation time. A similar idea has been proposed in [3], where $(dd|dd)$ is partitioned in $i$, $j$, and $k$ indices. $(ff|ff)$ is partitioned in $i$ and $j$ indices.
Algorithm 1 1qnt multi-level reduction algorithm (without 8-fold symmetry)
0: Input value shell quartet id, basis info
0: Output value $J$ and $K$ matrices
[…]
2: //Distribute fragments to different threads
3:for $f_{i}=0$ to $N^{f}_{i}/F_{i}$ in parallel do
4:  for $f_{j}=0$ to $N^{f}_{j}/F_{j}$ in parallel do
5:   for $f_{k}=0$ to $N^{f}_{k}/F_{k}$ in parallel do
6:    for $f_{l}=0$ to $N^{f}_{l}/F_{l}$ in parallel do
7:     // Integral fragment $T_{ijkl}=(ij|kl)$
8:     $T^{\text{reg}}[\colon N^{f}_{i},\colon N^{f}_{j},\colon N^{f}_{k},\colon N^{f}_{l}]=0$
9:     for $i_{p}=0$ to $n_{i}^{p}$ do
10:      for $j_{p}=0$ to $n_{j}^{p}$ do
11:       for $k_{p}=0$ to $n_{k}^{p}$ do
12:        for $l_{p}=0$ to $n_{l}^{p}$ do
13:         // Compute integrals with Eq. (1)-(3)
14:         $T^{\text{reg}}[i,j,k,l]\mathrel{+}=c_{p}c_{q}c_{r}c_{s}(pq|rs)$
15:        end for
16:       end for
17:      end for
18:     end for
19:     Read $D_{J}^{\text{reg}}[:F_{k},:F_{l}]$ and $D_{K}^{\text{reg}}[:F_{j},:F_{l}]$
20:     $J^{\text{reg}}[\colon F_{i},\colon F_{j}]=0,K^{\text{reg}}[\colon F_{i},\colon F_{k}]=0$
21:     for $i=0$ to $F_{i}$ do
22:      for $j=0$ to $F_{j}$ do
23:       for $k=0$ to $F_{k}$ do
24:        for $l=0$ to $F_{l}$ do
25:         //Thread level reduction
26:         $J^{reg}[i,j]\mathrel{+}=\sum_{kl}T^{\text{reg}}[i,j,k,l]D_{J}^{\text{reg}}[k,l]$
27:         $K^{reg}[i,k]\mathrel{+}=\sum_{jl}T^{\text{reg}}[i,j,k,l]D_{K}^{\text{reg}}[j,l]$
28:        end for
29:       end for
30:      end for
31:     end for
32:     //Block level reduction
33:     $J^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,jF_{j}\!:\!(j+1)F_{j}]\mathrel{+}=J^{\text{reg}}[\!:\!F_{i},\!:\!F_{j}]$
34:     $K^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,kF_{k}\!:\!(k+1)F_{k}]\mathrel{+}=K^{\text{reg}}[\!:\!F_{i},\!:\!F_{k}]$
35:    end for
36:   end for
37:  end for
38:end for
39: atomicAdd($J[:\!N^{f}_{i},:\!N^{f}_{j}],J^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{j}]$)
40: atomicAdd($K[:\!N^{f}_{i},:\!N^{f}_{k}],K^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{k}]$)
Optimize the fragmentation scheme. The performance of the integral kernel is sensitive to the fragment size. Different fragmentation strategies can easily lead to a 10x difference in performance. The fragment size can impact the register usage of each thread, the shared memory usage of each thread block, and the number of shell quartet assigned to each block. The resource consumption also need to meet the hard requirements of CUDA compute capability. We have to take these factors into account for searching the optimal fragment size. This optimization problem is less likely to be solved analytically. We look for the optimal fragmentation strategy by grid search. The searching space of high angular momentum can be huge. We made several assumptions:
  • We assume that each dimension of the integral is a multiple of the fragment size. There probably exists a better tiling strategy without this constraint. However, based on the grid search results, we found that the divisible fragment size, in general, is optimal. An additional benefit of this constraint is that each thread does not need to check the bound of iteration, which saves the significant computational cost.
  • The optimal fragment size is independent of the number of primitives. This assumption in general is not true, since the kernel with more number of primitives take more registers. Then, less registers are left for the integral fragment. The assumption is made for compromising with the high computational cost. We search the optimal fragment size based on a fixed number of primitives for low-angular orbital (s, p orbitals). The number of primitives for high-angular atomic orbital (d, f, g orbitals) is set to be 1.
  • The block size is always 256 threads.
Since the double-precision and single-precision algorithms differ in their usage of shared memory, register count, and hardware units, a strategy optimized for double precision is not necessarily optimal for single precision. Therefore, we optimize fragmentation strategies separately for each precision. The default fragmentation schemes are tuned on NVIDIA A100-80G and A10-24G GPUs, respectively. To ensure compatibility, the fragment size is optimized under the constraint of 48 KB shared memory per streaming multiprocessor, allowing the strategy to be applicable to older GPU generations. An optimization script is included in the JoltQC release, enabling users to retune the fragment size for different GPU architectures if they diverge significantly from the default targets.
We benchmark the 1tnq algorithm with a straight peptide chain composed of 30 glycines (gly30) and def2-TZVPP [31] basis set. Density matrix screening is turned off. Results on NVIDIA A100-80G and A10-24G GPUs appear in Figure 4 and 5, respectively. Using the default fragmentation for FP64, the 1tnq algorithm is 5-8x faster than GPU4PySCF v1.4 on A100-80G. When switched to FP32, the algorithm is further accelerated by 2x. For high-angular-momentum integrals, the computational cost of exp and erf functions is not significant relative to multiply/add operations. On A10-24G, the high-angular-momentum integrals are accelerated by roughly 2x for FP64. This is due to the algorithm in GPU4PySCF v1.4 being memory bound. The 1qnt algorithm reduces the memory bottleneck by multi-level reduction. When switched to FP32, the integrals are accelerated by 10-20x. Further performance analysis is provided in Appendix (Section 8).
Figure 4: Averaged kernel runtime of JoltQC v0.1 JK calculation using 1qnt algorithm, on the gly30 system, def2-TZVPP basis set. One Nvidia A100-80G GPU is used. The performance is compared with GPU4PySCF v1.4. Top 12 most time-consuming kernels are selected, and are labeled as (angular momentum combination)/(contraction pattern combination).
Figure 5: Averaged kernel runtime of JoltQC v0.1 JK calculation using 1qnt algorithm, on the gly30 system, def2-TZVPP basis set. One Nvidia A10-24G GPU is used. The performance is compared with GPU4PySCF v1.4. Top 12 most time-consuming kernels are selected, and are labeled as (angular momentum combination)/(contraction pattern combination).
4 Results
4.1 Compilation Overhead
Just-in-time compilation introduces non-negligible overhead, but this cost is amortized through on-disk caching and in-memory reuse. When a kernel is first invoked, it is compiled and the resulting binary is saved to disk. The cached binaries can be reused across the process. During the first SCF iteration, each process loads these binaries into RAM directly; all subsequent iterations then fetch them directly from memory. Table 3 summarizes the one-time compilation times for various kernels using CUDA v12.1. As expected, high–angular-momentum integrals take longer to compile and generate larger binaries: for the C, H, O, and N elements with the 6-31G* basis (140 kernels), the compilation time is just 30 s and the total kernel size is roughly 10 MB, while the def2-TZVPP basis (750 kernels) yields about 100 MB of binary and takes approximately 200 s to compile. Once those kernels are compiled, subsequent calls take less than 1 s to load cached binaries.
These figures serve as a reference. Actual compilation times will vary with different CUDA versions and hardware quality, particularly disk performance. On modern devices, the compilation overhead can be reduced in various ways. In practice, especially in homogeneous cloud environments, it can be advantageous to store compiled kernels on shared or remote storage and download them as needed. The similar approach has been used in the machine learning community, e.g. PyTorch AOTInductor. Alternatively, retaining intermediate .ptx files and loading them directly allows users to bypass one step of compilation.
Integral pattern Algorithm
Compilation
time (ms)
Binary size
(ss|ss)/(11|11) 1Q1T 284.8 25
(ss|ss)/(33|33) 1Q1T 512.0 74
(pp|pp)/(11|11) 1Q1T 638.8 75
(pp|pp)/(33|33) 1Q1T 880.6 94
(dd|dd)/(11|11) 1QnT 6,011 194
(ff|ff)/(11|11) 1QnT 7,697 270
(gg|gg)/(11|11) 1QnT 59,468 602
Table 3: Compilation cost of selected integral kernels with NVRTC in CUDA V12.1. All of the kernels are compiled in FP64.
4.2 Benchmark with GPU4PySCF and TeraChem
In this work, integrals for the Coulomb (J) and exchange (K) terms are screened simultaneously. GPU4PySCF and JoltQC use the same screening technique. The number of significant shell quartets in each JK calculation is identical in both cases. In double-precision, JoltQC reproduces GPU4PySCF’s results exactly, so we omit a detailed FP64 accuracy analysis. Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
We use a customized TeraChem v1.9 build with f-orbital support. Although both GPU4PySCF and TeraChem employ the Schwarz inequality to bound integral magnitudes, they apply it at different levels: TeraChem at the primitive shell level and GPU4PySCF at the contracted level. Their SCF initial guesses and DIIS settings also differ, so our GPU4PySCF–TeraChem comparisons are not strictly apples-to-apples. For all calculations shown, the screening threshold for both J and K matrices is set to $10^{-13}$. SCF with GPU4PySCF/JoltQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations. This does not imply TeraChem’s settings are superior, as convergence behavior can vary with the system. A sample TeraChem input is provided in Appendix 7. All the basis sets are in Cartesian format.
When both codes run in FP64, their total energies agree to within 0.001 mHa. When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both JoltQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations. A more practical mixed-precision algorithm is among our future directions.
Molecule Basis
GPU4PySCF/FP64
- JoltQC/FP32
GPU4PySCF/FP64
- TeraChem/FP64
JoltQC/FP32
- TeraChem/FP32
gly30 6-31G* 0.23 mHa 0.0005 mHa -0.07 mHa
gly30 def2-TZVPP 1.95 mHa 0.0004 mHa 2.10 mHa
valinomycin 6-31G* -0.06 mHa -0.001 mHa -0.26 mHa
valinomycin def2-TZVPP 0.68 mHa -0.001 mHa 0.63 mHa
Table 4: The accuracy of Hartree-Fock (HF) energy with JoltQC v0.1 against TeraChem v1.9. GPU4PySCF/FP64 and JoltQC/FP64 matches exactly and is not shown in the table.
As shown in Table 5, JoltQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis. Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal. In double-precision, JoltQC outperforms TeraChem by about 4x when def2-TZVPP basis set is used, and a less significant 1.6x for 6-31G* basis. In single-precision, JoltQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP. Both TeraChem and JoltQC still employ double precision for eigenvalue decomposition and DIIS, making these linear-algebra routines a significant cost on the A10-24 GB GPU. Isolating the JK kernel, JoltQC is approximately three times faster than TeraChem.
Molecule Protocol GPU4PySCF TeraChem JoltQC
JK/cycle SCF JK/cycle SCF JK/cycle SCF
gly30 HF/6-31G*/A100/FP64 1.9 s 28.1 s 1.3 s 19.1 s 0.8 s 10.4 s
gly30 HF/def2-TZVPP/A100/FP64 78.4 s 1,125.7 s 91.4 s 1,055.1 s 18.8 s 268.1 s
gly30 HF/6-31G*/A10/FP32 - - 1.0 s 23.1 s 0.7 s 16.5 s
gly30 HF/def2-TZVPP/A10/FP32 - - 63.5 s 870.8 s 20.0 s 406.1 s
Table 5: Benchmark among GPU4PySCF v1.4, TeraChem v1.9, and JoltQC v0.1, on A100-80G using FP64 arithmetic, and on A10-24G using FP32 arithmetic.
5 Conclusion
We optimized the low–angular-momentum electron repulsion integral algorithm in GPU4PySCF v1.4 using just-in-time (JIT) compilation and introduced a novel high–angular-momentum integral method. The core CUDA implementation for both algorithms spans only 1,000 lines of code in total. JIT compilation not only accelerates the evaluation of Gaussian-type-orbital integrals but also provides developers with greater flexibility. Our streamlined implementation delivers a 2-4x performance improvement over GPU4PySCF v1.4 and is significantly easier to maintain. Additionally, it supports single- and double-precision arithmetic, accepts multiple density matrices as input, and integrates seamlessly with the existing GPU4PySCF framework.
6 Future Work and Limitations
High level abstractions of JIT. Although our current implementation relies heavily on handwritten CUDA kernels, it would be preferable to express these routines in a high-level language such as Triton [28], Numba [13], or JAX [5]. These frameworks offer richer support for kernel fusion, making it easier to embed quantum-chemistry operations within machine-learning pipelines. For example, one could fuse the on-the-fly computation of integrals with subsequent matrix–matrix multiplications (not only contraction with density matrices), yielding more efficient end-to-end workflows. Further, one could write a customized kernel for the derivatives with respect to contraction coefficients and exponents, then port to an existing auto-differentiation framework [34].
Supporting more functionalities. A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels. The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post Hartree-Fock methods, and a wide array of molecular properties are all essential. By integrating JoltQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
Mixed precision algorithms. This paper focuses on the performance of the double precision and single precision separately. Thanks to the Schwartz inequality, the computational load of the integrals can be naturally split into FP64 calculations and FP32 calculations without loss of accuracy theoretically. Other research efforts [17, 20] show that 90% of the computations can be in FP32. A naive mixed precision strategy can be easily proposed based on this work. One should notice, even though the value of integral can be represented in FP32, not every underlying arithmetic operation for computing the integral can necessarily be computed with FP32. More effort is needed to rigorously implement a strategy without error accumulation.
Screening techniques. The intrinsic scalings of the Coulomb and exchange terms are different. The number of significant Coulomb terms increases quadratically, while the number of significant exchange terms increases linearly with respect to the system size. It is more efficient to evaluate the Coulomb and exchange terms separately. In some quantum chemistry packages such as TeraChem, the Coulomb terms are efficiently evaluated with J-engine [32]. Besides the Schwartz inequality, which gives a rigorous upper bound, other estimates can further eliminate the unnecessary calculations. For example, the MBIE [14], QQR [18], and CSAM [27]. MBIE provides a rigorous upperbound, but it is expensive to calcuate. QQR and CSAM take advantage of distance dependence in their estimates, hence improving the performance for large systems. But those estimates are less rigorous.
Author Contributions
Xiaojie Wu led the conceptualization of JIT algorithms and developed JoltQC v0.1, and wrote the initial draft. Qiming Sun contributed the improvements of the JK kernels v1.0 to v1.4 in GPU4PySCF, which are not documented elsewhere. Yuanheng Wang explored single-precision algorithms in GPU4PySCF, discussed JIT algorithms in JoltQC v0.1, and wrote the manuscript.
Acknowledgments
The authors appreciate the support of Dr. Wen Yan and the ByteDance Seed management team. Special thanks are also extended to Dr. Zhichen Pu for his discussions. The authors also thank the PySCF community for providing valuable tools that contributed to this work, especially the GPU4PySCF contributors.

### v6

@@
 Algorithms in electronic structure theory often require handling a combinatorial variety of tensor contraction patterns and sparsity conditions that depend on the input molecular system.
-Static compilation must conservatively support all conditional branches, hyperparameter combinations, and numerical thresholds, leading to bloated binaries, excessive control flow divergence, poor cache and register utilization, and inefficient use of memory bandwidth.
+Static compilation must conservatively support all conditional branches, parameter combinations, and numerical thresholds, leading to bloated binaries, excessive control flow divergence, poor cache and register utilization, and inefficient use of memory bandwidth.
 On GPUs, these inefficiencies are further magnified by strict constraints on register pressure, warp divergence, and hardware occupancy, often requiring significant manual optimization.
-For example, most quantum chemistry packages deal with low-angular-momentum integrals successfully but still struggle with the performance of high-angular momentum integrals, which suffer from all of these problems.
+For example, most quantum chemistry packages deal with low-angular-momentum integrals successfully but still struggle with the performance of high-angular-momentum integrals, which suffer from all of these problems.
 Just-in-time (JIT) compilation offers a transformative solution by enabling runtime code specialization based on the actual inputs: e.g., integral angular-momentum pattern, the number of primitives in an orbital shell, and task type.
@@
 This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures.
-Furthermore, the developers are able to design novel algorithms which are not general enough under the AOT framework.
+Furthermore, developers can explore specialized algorithms that would be impractical to hand-code under the AOT model.
 In this paper, we demonstrate several algorithms due to this benefit.
@@
 These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
-JIT compilation benefits mixed precision algorithms in quantum chemistry, which naturally accelerate electron repulsion integrals by exploiting the Schwarz inequality [17, 20].
-Most tasks can be calculated in single-precision (FP32) as compared to the commonly-used double-precision (FP64).
+The Schwarz inequality provides rigorous upper bounds on integral magnitudes, enabling most shell quartets to be evaluated in single precision (FP32) rather than the conventional double precision (FP64) [18, 21].
+JIT compilation makes this mixed-precision approach particularly easy to implement: since precision is a compile-time parameter, the same source code generates both FP32 and FP64 kernels without code duplication.
+Most integral tasks can be calculated in single-precision as compared to the commonly-used double-precision.
 Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads.
-These benefits can be better utilized by JIT compilation without significant changes to the code.
+JIT compilation exploits all three advantages automatically: the same kernel source is compiled once for FP32 and once for FP64, with the compiler selecting the appropriate hardware intrinsics and register allocation for each precision — no manual code duplication required.
 Contributions: 1) We introduce JIT to the quantum chemistry community and optimize the two-electron integral kernels (1q1t in Sec.
+3.1 and 1qnt in Sec.
 3.2) with compilation techniques.
@@
 Since each of the Gaussian type orbitals is contracted from a set of primitive Gaussian functions [12], the integral tensor for a shell quartet is expressed as the sum of the primitive integral tensors
-$(ij|kl)=\sum_{pqrs}c_{p}c_{q}c_{r}c_{s}(pq|rs)$ (1)
-$c_{p}$, $c_{q}$}, $c_{r}$, and $c_{s}$ are the pre-defined contraction coefficients for each shell in the basis set.
-The number of Gaussian functions in each primitive shell is $l(l+1)/2$ or $2l+1$ for cartesian atomic orbitals or spherical atomic orbitals respectively, where $l$ is the angular momentum of the shell.
-Only Cartesian orbitals are considered in the following sections, and spherical orbitals are just a linear combination of Cartesian orbitals with fixed transformation coefficients.
+$(ij|kl)=\sum_{p\in i}\sum_{q\in j}\sum_{r\in k}\sum_{s\in l}c_{ip}\,c_{jq}\,c_{kr}\,c_{ls}\;(pq|rs).$ (1)
+$c_{ip}$, $c_{jq}$, $c_{kr}$, and $c_{ls}$ are the pre-defined contraction coefficients for each shell in the basis set.
+The number of Gaussian functions in each primitive shell is $(l+1)(l+2)/2$ or $2l+1$ for cartesian atomic orbitals or spherical atomic orbitals respectively, where $l$ is the angular momentum of the shell.
+Only Cartesian orbitals are considered in the following sections, and spherical orbitals are just linear combinations of Cartesian orbitals with fixed transformation coefficients.
 These linear transformations can be performed efficiently outside the kernel.
-As a result, each dimension of the primitive integral $(pq|rs)$ is $N^{f}=l(l+1)/2$.
+As a result, each dimension of the primitive integral $(pq|rs)$ is $N^{f}=(l+1)(l+2)/2$.
 We denote the dimension of the integral for each shell quartet as $(N_{i}^{f},N_{j}^{f},N_{k}^{f},N_{l}^{f})$.
@@
 $(pq|rs)=\sum_{r}^{N_{r}}(pq|rs)(t_{r})\cdot w_{r}\\$ (2)
-where $N_{r}$ is the number of Rys roots.
+where $N_{\text{r}}=\left\lfloor(l_{i}+l_{j}+l_{k}+l_{l})/2\right\rfloor+1$ is the number of Rys roots.
 $t_{r}$ and $w_{r}$ are the roots and weights of Rys quadrature.
@@
 $I_{x}$, $I_{y}$, and $I_{z}$ are the intermediate variables introduced in [7, 9].
-Each of them is a tensor of shape $(l_{i}+1,l_{j}+1,l_{k}+1,l_{l}+1)$.
-They depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only.
+For every root, each of the intermediate variables is a tensor of shape $(l_{i}+1,l_{j}+1,l_{k}+1,l_{l}+1)$.
+The total number of intermediate variables is $3\times 3\times(2\times 2\times 2\times 2)=144$ for calculating a $(pp|pp)$ integral (the first 3 refers to the three Cartesian dimensions, and the second 3 is the number of roots).
+For $(gg|gg)$ integral, the number of variables is raised up to $3\times 9\times(5\times 5\times 5\times 5)=16875$, which exceeds the capacity of GPU registers.
+The size of intermediate variables can be reduced if the roots are handled sequentially.
+The same space can be reused for different roots.
+The intermediate variables $I_{x},I_{y},\text{and }I_{z}$ depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only.
 They can be efficiently evaluated using recursive relations, documented in [9].
@@
 3.2.
-The index set for enumerating the Guassian functions of a primitive shell is defined as
+The index set for enumerating the Gaussian functions of a primitive shell is defined as
 $S_{l}:=\{(i,a,b,c)|a+b+c=l,i=\texttt{idx}(a,b,c)\}$
 The idx function is defined differently in various quantum chemistry packages.
-One will need to double check the conventions to align the quantities in atomic basis, such as density matrix, molecular orbitals, and so on.
-We adopt the convention of reversed lexicographic order.
+We adopt reversed lexicographic order: for $l=2$ (d shell), $\texttt{idx}(2,0,0)=0$, $\texttt{idx}(1,1,0)=1$, …, $\texttt{idx}(0,0,2)=5$.
+Note that other packages may adopt different orderings; the density matrix and orbital coefficients must be reordered accordingly.
 The indexing operation in general makes the memory access random and less efficient.
-Because of its huge size, the four-dimensional integral tensor is not formed explicitly in the main storage, but instead contracted with the density matrix $D$ to form the Coulomb ($J$) and exchange ($K$) potential matrices:
+Because of its huge size, the four-dimensional integral tensor is not formed explicitly in the main memory, but instead contracted with the density matrix $D$ to form the Coulomb ($J$) and exchange ($K$) potential matrices:
 $J_{ij}=\sum_{kl}(ij|kl)D_{kl},\text{ and }K_{ik}=\sum_{jl}(ij|kl)D_{jl}$ (4)
-For a molecular system with real-valued orbitals, the 8-fold symmetry of $(ij|kl)$ can be used to save computational cost.
-In particular, for the integral element $(ij|kl)$ composed of the $i$-th, $j$-th, $k$-th and $l$-th atomic orbital, it is contracted with two density matrix elements $D_{kl}$ and $D_{lk}$ to form two matrix elements $J_{ij}$ and $J_{ji}$, and it is contracted with four density matrix elements $D_{jl}$, $D_{jk}$.
+For a molecular system with real-valued orbitals, the 8-fold symmetry of $(ij|kl)$ can be used to save computational cost,
+(ij|kl)=(ji|kl)=(ij|lk)=(ji|lk)=(kl|ij)=(lk|ij)=(kl|ji)=(lk|ji).(ij|kl)=(ji|kl)=(ij|lk)=(ji|lk)=(kl|ij)=(lk|ij)=(kl|ji)=(lk|ji). (5)
+As a result, only integrals $(ij|kl)$ where $i>j$, $k>l$ and $iN_{ao}+j>kN_{ao}+l$ needs to be evaluated, and this saves $7/8$ of the integral evaluation work.
+In particular, for the integral element $(ij|kl)$ composed of the $i$-th, $j$-th, $k$-th and $l$-th atomic orbital, the J matrix requires a trace over the bra-ket pair $(k,l)$, summing $(ij|kl)D_{kl}$, yielding two updates ($J_{ij}$ and $J_{ji}$ by symmetry).
+The K matrix involves cross-contractions where one index from the bra pairs and one from the ket, yielding four distinct updates.
+Therefore, the integral $(ij|kl)$ is contracted with two density matrix elements $D_{kl}$ and $D_{lk}$ to form two matrix elements $J_{ij}$ and $J_{ji}$.
+And it is simultaneously contracted with four density matrix elements $D_{jl}$, $D_{jk}$.
 $D_{il}$, $D_{ik}$ to form four $K$ matrix elements $K_{ik}$, $K_{il}$, $K_{jk}$, $K_{jl}$.
 Conventionally, the Schwarz inequality for shell quartets
-$|(ij|kl)|\leq\max_{i\in S_{l_{i}},j\in S_{l_{j}}}\sqrt{(ij|ij)}\cdot\max_{k\in S_{l_{k}},l\in S_{l_{l}}}\sqrt{(kl|kl)}$ (5)
+$|(ij|kl)|\;\leq\;\sqrt{(ij|ij)}\sqrt{(kl|kl)}.$ (6)
 is used to screen the significant shell quartets, i.e.
@@
 incremental SCF), respectively.
-Since $\Delta D$ is getting smaller along the SCF iterations, the screening is more effective.
+Since $\Delta D$ decreases along the SCF iterations, the screening is more effective.
 In practice, the density matrix screening is performed at shell pair level, which is similar to Eq.
@@
 3 Design and Algorithms
-The input of a quantum chemistry calculation contains the protocol of the method, geometry of a molecule, and various thresholds.
-These inputs are unknown when a typical quantum chemistry package is compiled.
-Therefore, all types of input have to be handled as dynamic variables.
-The idea of JIT design is to feed the input information to the compiler at compilation time such that the compiler can aggressively optimize the code.
-On the other hand, if the compilation completely depends on all the input information, i.e.
-if we recompile for every job even if successively jobs are only different in geometry, the recompilation may slow down the overall execution time, since the compilation can be more expensive than the actual calculation with an improper design.
-In the design of JIT programs, distinguishing between static and dynamic variables is crucial for user experience and performance.
-In general, we follow this principle: The information of quantum chemistry protocol and device information are static for compilation.
-That is, when the quantum chemistry method, basis set, or type of GPUs is changed, the kernels should be recompiled.
-The same kernels will be used for different molecules.
-All the molecule information, including the atomic numbers, atomic types and atomic coordinates, and the total charge and spin of the system, are dynamic variables for compilation.
+The central design decision in a JIT framework is which inputs to treat as compile-time constants (static) and which to pass as runtime arguments (dynamic).
+Making too few variables static forfeits optimization opportunities; making too many static causes excessive recompilation.
+We adopt the following rule: protocol and device information — the quantum chemistry method, basis set type, and GPU architecture — are static, while molecular information — atomic coordinates, Gaussian orbital exponents, and contraction coefficients — remains dynamic.
+This ensures that a single compilation serves all molecules sharing the same basis set, while still enabling aggressive loop unrolling and register allocation.
 More specifically, we select the angular momentum and contraction pattern of the basis set, as well as the task type, as the template parameters for the kernels.
 They are described in Table 1.
-Once these variables are fixed, the size of any intermediate variable in the integral kernels is known at compilation time.
+A crucial consequence of this design: once the angular momentum, contraction pattern, and number of primitives are fixed, every loop bound and array size in the integral kernel becomes a compile-time constant.
+The compiler can then fully unroll loops, allocate registers statically, and eliminate all branching.
 If those parameters are changed, we have to recompile the kernels.
-With such choices, we avoid complicated runtime logic flows for different data type, task type, and unrolling strategies.
+With such choices, we avoid complicated runtime logic flows for different data types, task types, and unrolling strategies.
 Thus the amount of source code is minimal: only ~1,000 lines of code for each algorithm, much less than the ~20,000 lines in GPU4PySCF v1.4.
-Template
-Parameter
-Data Type Description
+Template Parameter Data Type Description
 $(l_{i},l_{j},l_{k},l_{l})$ int Angular momentum of shells
@@
 Since many elements share the same contraction pattern but different coefficients and exponents, we do not need to compile multiple kernels for the same contraction pattern.
-Generally, reading those data is not the bottleneck of the kernels.
+The basis set parameters (exponents, coefficients, coordinates) are stored in global memory and read once per shell quartet.
+Since the data size is usually small and is efficiently cached through the L1/L2 cache, reading those data is not the bottleneck of the kernels.
+Another strategy is to form the basis pair data as the kernel input, and the pair data can be more efficient in exp operation in some scenarios, for example F64 kernel on RTX cards, at the cost of less efficient cache utilization.
+Directly taking the basis data as input is part of the current design, but may not be optimal.
 All of the dynamic parameters are summarized in 2.
-Input
-Data Type Description
+Input Data Type Description
 $(c_{i},c_{j},c_{k},c_{l})$ double/single contraction coefficients in basis
@@
 Figure 1: Workflow of SCF iterations with NVRTC as JIT compiler.
+Based on this design, we build heuristics (including algorithms and fragmentation strategies) to deliver optimized kernels for a specific device type and precision type (FP64/FP32).
+The heuristics are benchmarked ahead of time, and the best ones are saved in a database as a part of the final package.
+In the runtime, the kernel caller will find the corresponding algorithm based on the device type and required precision.
+Then, only one optimized kernel is compiled.
+It should be noted that, even for the exact same integrals, with different device/precision combination, different heuristics might be selected.
+In particular, the FP64 algorithm may be different from the FP32 algorithm for the same angular momentum, in order to achieve the best performance.
+In the next two sections, we present two heuristics algorithms for integral kernels, each targeting a different angular-momentum regime.
+The 1q1t algorithm (Sec.
+3.1) assigns one thread per shell quartet and is often well-suited for low angular momentum (s, p shells), where the integrals fit in registers.
+The 1qnt algorithm (Sec.
+3.2) fragments the integral across multiple threads and is designed for high angular momentum (d, f, g shells), where register pressure becomes the bottleneck.
 3.1 One quartet one thread (1q1t) algorithm
-The first algorithm we present is largely the same as the classical Rys quadrature algorithm on CPUs [26], except for a more aggressive loop unrolling strategy is used.
+The first algorithm we present is largely the same as the classical Rys quadrature algorithm on CPUs [27], except that a more aggressive loop unrolling strategy is used.
 First, a list of significant shell quartets is generated by running over all possible combinations screened by the Schwartz inequality.
@@
 (1) (including all primitives in a contraction and all functions in a shell), and contract it with the density matrix according to Eq.
-(4), until the pool of all significant shell quartet is empty.
+(4).
 This algorithm was implemented in the early version of GPU4PySCF (v1.0 and before).
 This straightforward implementation has two performance issues.
-1) There are still multiple loops not unrolled due to unknown variables at compilation time, such as number of primitivies for each shell and number of density matrices.
-This is actually the main bottleneck of low-angular-momentum integrals.
-2) Significant register spill out onto local memory for high angular momentum.
-In this subsection, we will show that with more known variables at compilation time, the performance is significantly improved.
+1) Several performance-critical loops — over primitives in each shell and over density matrices — have bounds unknown at compile time and therefore cannot be unrolled.
+On GPUs, unrolled loops enable the compiler to schedule instructions more efficiently, eliminate branch overhead, and allocate registers to specific loop iterations.
+This is the dominant bottleneck for low-angular-momentum integrals, which are individually cheap but numerous.
+JIT compilation resolves this by making all loop bounds compile-time constants.
+2) Significant register spills to local memory for high angular momentum.
+In this subsection, we show that with more known variables at compilation time, the performance is significantly improved.
 The second issue will be discussed in the next subsection.
@@
 In the single-precision implementation, the exponents, contraction coefficients, and atomic coordinates are prepared and loaded in FP32.
-The density matrix values are loaded as FP64 and then casted to FP32 in kernels.
+The density matrix is maintained in FP64 throughout the SCF procedure for diagonalization and DIIS, so we load it in its native format and cast to FP32 inside the kernel to avoid maintaining a separate single-precision copy.
+One may improve the memory throughput by maintaining a separate single-precision copy.
 Inside the kernel, all the arithmetic operations are performed in FP32 and accelerated with the --use_fast_math compiler flag.
-One should notice that fast-math intrinsics for single precision is much faster, but not IEEE-754 accurate.
-When $J$ and $K$ are calculated, they are first casted to FP64, then accumulated with atomicAdd in FP64.
+The --use_fast_math flag replaces standard math functions (sqrt, exp, erf) with hardware-accelerated intrinsics that are faster but not IEEE-754 compliant (i.e.
+they introduce an additional small rounding error).
+The accelerated exp function benefits the evaluations of Gaussian functions, as the kernel takes the native Gaussian exponents and contraction coefficients as inputs.
+The pair data is calculated inside the kernel.
+The benefit of exp is negligible if one precompute the pair data before invoking the JK kernel.
+The calculations of Rys roots and weights are accelerated as well for two reasons.
+1) We directly use erf to evaluate the root and weight when $N_{r}=1$ (interpolation is used for other cases), as we found that directly evaluating the special functions is often faster than the interpolation.
+2) In the long-distance regime for $N_{r}>1$ (outside of interpolation window) of Rys roots and weights, sqrt is used in the asymptotic formula.
+[13]
 We benchmark the performance of JK kernels with a straight peptide chain composed of 120 glycines (gly120), and 6-31G [6] basis set.
@@
 Thanks to the special hardware acceleration of these functions on GPU, the overall speedup of FP32 reaches about 5-10x on A100-80G.
-Since A10-24G has 1/32 hardware units for FP64 compared to FP32, FP64 algorithms are always bounded by the limited compute units.
+Since the A10-24G only delivers 1/32 of P64 throughput compared to FP32, FP64 algorithms are always bounded by the limited compute units.
 The improvement with JIT is marginal.
@@
 3.2 Quartet fragmentation (1qnt) algorithm
-For high-angular momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers.
-From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy is introduced to split the integral into multiple threads.
-Instead of one block thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet.
-The strategy significantly improved the performance of the high-angular integrals.
+For high-angular-momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers.
+From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy was introduced to split the integral into multiple threads.
+Instead of a single thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet.
+The strategy significantly improved the performance of the high-angular-momentum integrals.
 In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory).
-Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variable $I_{x}$, $I_{y}$ and $I_{z}$ will be stored in shared memory.
+Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variables $I_{x}$, $I_{y}$ and $I_{z}$ are stored in shared memory.
 Then, each thread in the group will compute and store into registers only a fraction of the integrals.
-For example, the shape of the integral (dd|dd) is (6,6,6,6), each thread computes a fragment with the size of (3,3,3,3).
-16 threads are needed to store the entire (dd|dd) integral.
-In the reduction stage, each thread is responsible for a fragment of 3x3 $v_{j}$ and $v_{k}$.
-The fragments $v_{j}$ and $v_{k}$ are further reduced at the thread block level.
+For example, for $(dd|dd)$, the integral tensor has shape (6,6,6,6).
+With fragment size (3,3,3,3), each dimension is split into 6/3 = 2 parts.
+The four fragment indices ($f_{i}$, $f_{j}$, $f_{k}$, $f_{l}$) are distributed across threads, requiring $2^{4}=16$ threads per shell quartet.
+After computing its integral fragment $T^{reg}$, each thread contracts $T^{reg}$ with the density matrix to produce local contributions to J and K.
+We denote the thread-local J contribution as $v_{j}$ (shape $F_{i}\times F_{j}$) and the thread-local K contribution as $v_{k}$ (shape $F_{i}\times F_{k}$).
+These are accumulated in registers during the thread-level reduction (Algorithm 1, lines 21–30), then reduced across threads at the block level (lines 32–36).
 The reduction scheme is presented in Algorithm 1.
 The shape of the intermediate variables are also known at compilation time.
-A similar idea has been proposed in [3], where $(dd|dd)$ is partitioned in $i$, $j$, and $k$ indices.
-$(ff|ff)$ is partitioned in $i$ and $j$ indices.
+A similar idea has been proposed in [3], where $(dd|dd)$ is partitioned in $i$, $j$, and $k$ indices, while $(ff|ff)$ is partitioned in $i$ and $j$ indices.
 Algorithm 1 1qnt multi-level reduction algorithm (without 8-fold symmetry)
@@
 0: Output value $J$ and $K$ matrices
+1: // Superscripts: reg = per-thread registers, shd = per-block shared memory, no superscript = global memory.
 […]
-2: //Distribute fragments to different threads
-3:for $f_{i}=0$ to $N^{f}_{i}/F_{i}$ in parallel do
-4:  for $f_{j}=0$ to $N^{f}_{j}/F_{j}$ in parallel do
-5:   for $f_{k}=0$ to $N^{f}_{k}/F_{k}$ in parallel do
-6:    for $f_{l}=0$ to $N^{f}_{l}/F_{l}$ in parallel do
-7:     // Integral fragment $T_{ijkl}=(ij|kl)$
-8:     $T^{\text{reg}}[\colon N^{f}_{i},\colon N^{f}_{j},\colon N^{f}_{k},\colon N^{f}_{l}]=0$
-9:     for $i_{p}=0$ to $n_{i}^{p}$ do
-10:      for $j_{p}=0$ to $n_{j}^{p}$ do
-11:       for $k_{p}=0$ to $n_{k}^{p}$ do
-12:        for $l_{p}=0$ to $n_{l}^{p}$ do
-13:         // Compute integrals with Eq. (1)-(3)
-14:         $T^{\text{reg}}[i,j,k,l]\mathrel{+}=c_{p}c_{q}c_{r}c_{s}(pq|rs)$
-15:        end for
+3: //Distribute fragments to different threads
+4:for $f_{i}=0$ to $N^{f}_{i}/F_{i}$ in parallel do
+5:  for $f_{j}=0$ to $N^{f}_{j}/F_{j}$ in parallel do
+6:   for $f_{k}=0$ to $N^{f}_{k}/F_{k}$ in parallel do
+7:    for $f_{l}=0$ to $N^{f}_{l}/F_{l}$ in parallel do
+8:     // Integral fragment $T_{ijkl}=(ij|kl)$
+9:     $T^{\text{reg}}[\colon\!N^{f}_{i},\colon\!N^{f}_{j},\colon\!N^{f}_{k},\colon\!N^{f}_{l}]=0$ // Allocate integral fragment in registers
+10:     for $i_{p}=0$ to $n_{i}^{p}$ do
+11:      for $j_{p}=0$ to $n_{j}^{p}$ do
+12:       for $k_{p}=0$ to $n_{k}^{p}$ do
+13:        for $l_{p}=0$ to $n_{l}^{p}$ do
+14:         // Compute integrals with Eq. (1)-(3)
+15:         $T^{\text{reg}}[i,j,k,l]\mathrel{+}=c_{ip}c_{jq}c_{kr}c_{ls}(pq|rs)$
 16:        end for
@@
 18:      end for
-19:     Read $D_{J}^{\text{reg}}[:F_{k},:F_{l}]$ and $D_{K}^{\text{reg}}[:F_{j},:F_{l}]$
-20:     $J^{\text{reg}}[\colon F_{i},\colon F_{j}]=0,K^{\text{reg}}[\colon F_{i},\colon F_{k}]=0$
-21:     for $i=0$ to $F_{i}$ do
-22:      for $j=0$ to $F_{j}$ do
-23:       for $k=0$ to $F_{k}$ do
-24:        for $l=0$ to $F_{l}$ do
-25:         //Thread level reduction
-26:         $J^{reg}[i,j]\mathrel{+}=\sum_{kl}T^{\text{reg}}[i,j,k,l]D_{J}^{\text{reg}}[k,l]$
-27:         $K^{reg}[i,k]\mathrel{+}=\sum_{jl}T^{\text{reg}}[i,j,k,l]D_{K}^{\text{reg}}[j,l]$
-28:        end for
+19:     end for
+20:     Read $D_{J}^{\text{reg}}[:\!F_{k},:\!F_{l}]$ and $D_{K}^{\text{reg}}[:\!F_{j},:\!F_{l}]$ // Read density matrix subblocks in registers
+21:     $J^{\text{reg}}[\colon\!F_{i},\colon\!F_{j}]=0,K^{\text{reg}}[\colon\!F_{i},\colon\!F_{k}]=0$ // Allocate accumulators for J, K in registers
+22:     for $i=0$ to $F_{i}$ do
+23:      for $j=0$ to $F_{j}$ do
+24:       for $k=0$ to $F_{k}$ do
+25:        for $l=0$ to $F_{l}$ do
+26:         //Thread level reduction
+27:         $J^{reg}[i,j]\mathrel{+}=\sum_{kl}T^{\text{reg}}[i,j,k,l]D_{J}^{\text{reg}}[k,l]$
+28:         $K^{reg}[i,k]\mathrel{+}=\sum_{jl}T^{\text{reg}}[i,j,k,l]D_{K}^{\text{reg}}[j,l]$
 29:        end for
@@
 31:      end for
-32:     //Block level reduction
-33:     $J^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,jF_{j}\!:\!(j+1)F_{j}]\mathrel{+}=J^{\text{reg}}[\!:\!F_{i},\!:\!F_{j}]$
-34:     $K^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,kF_{k}\!:\!(k+1)F_{k}]\mathrel{+}=K^{\text{reg}}[\!:\!F_{i},\!:\!F_{k}]$
-35:    end for
+32:     end for
+33:     //Block level reduction
+34:     $J^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,jF_{j}\!:\!(j+1)F_{j}]\mathrel{+}=J^{\text{reg}}[\!:\!F_{i},\!:\!F_{j}]$
+35:     $K^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,kF_{k}\!:\!(k+1)F_{k}]\mathrel{+}=K^{\text{reg}}[\!:\!F_{i},\!:\!F_{k}]$
 36:    end for
@@
 38:  end for
-39: atomicAdd($J[:\!N^{f}_{i},:\!N^{f}_{j}],J^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{j}]$)
-40: atomicAdd($K[:\!N^{f}_{i},:\!N^{f}_{k}],K^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{k}]$)
+39:end for
+40: atomicAdd($J[:\!N^{f}_{i},:\!N^{f}_{j}],J^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{j}]$) // Write block results to global J
+41: atomicAdd($K[:\!N^{f}_{i},:\!N^{f}_{k}],K^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{k}]$) // Write block results to global K
 Optimize the fragmentation scheme.
@@
 We have to take these factors into account for searching the optimal fragment size.
-This optimization problem is less likely to be solved analytically.
+This optimization problem is unlikely to be solved analytically.
 We look for the optimal fragmentation strategy by grid search.
+For each candidate fragmentation, we compile and execute the kernel on a representative workload and select the configuration with the lowest measured runtime.
 The searching space of high angular momentum can be huge.
@@
 Then, less registers are left for the integral fragment.
-The assumption is made for compromising with the high computational cost.
-We search the optimal fragment size based on a fixed number of primitives for low-angular orbital (s, p orbitals).
-The number of primitives for high-angular atomic orbital (d, f, g orbitals) is set to be 1.
+This assumption is highly empirical.
+In actual runs, kernels with more primitives consume more registers, leaving fewer for the integral fragment.
+However, jointly optimizing over both fragment size and primitive count would make the search space intractable.
+We therefore fix the primitive count at a representative value and search only over the space of fragment size.
+For high-angular-momentum shells ($d$, $f$, $g$), which typically have fewer primitives, we set the primitive count to 1.
   • The block size is always 256 threads.
+We optimize fragmentation strategies separately for the combinations of precisions and GPU devices.
+As previously mentioned, the FP64 algorithm may use a different fragmentation scheme from the FP32 algorithm for the same angular momentum.
 Since the double-precision and single-precision algorithms differ in their usage of shared memory, register count, and hardware units, a strategy optimized for double precision is not necessarily optimal for single precision.
-Therefore, we optimize fragmentation strategies separately for each precision.
 The default fragmentation schemes are tuned on NVIDIA A100-80G and A10-24G GPUs, respectively.
@@
 An optimization script is included in the JoltQC release, enabling users to retune the fragment size for different GPU architectures if they diverge significantly from the default targets.
-We benchmark the 1tnq algorithm with a straight peptide chain composed of 30 glycines (gly30) and def2-TZVPP [31] basis set.
+We benchmark the 1qnt algorithm with a straight peptide chain composed of 30 glycines (gly30) and def2-TZVPP [32] basis set.
 Density matrix screening is turned off.
 Results on NVIDIA A100-80G and A10-24G GPUs appear in Figure 4 and 5, respectively.
-Using the default fragmentation for FP64, the 1tnq algorithm is 5-8x faster than GPU4PySCF v1.4 on A100-80G.
+Using the default fragmentation for FP64, the 1qnt algorithm is 5-8x faster than GPU4PySCF v1.4 on A100-80G.
 When switched to FP32, the algorithm is further accelerated by 2x.
@@
 4 Results
+We evaluate JoltQC along three axes: compilation overhead (Sec.
+4.1), end-to-end SCF performance against GPU4PySCF v1.4 and TeraChem v1.9 (Sec.
+4.2), and single-precision accuracy.
+All benchmarks use a single NVIDIA A100-80G GPU for FP64 and an A10-24G GPU for FP32 unless otherwise noted.
 4.1 Compilation Overhead
@@
 When a kernel is first invoked, it is compiled and the resulting binary is saved to disk.
-The cached binaries can be reused across the process.
+The cached binaries can be reused across the processes.
 During the first SCF iteration, each process loads these binaries into RAM directly; all subsequent iterations then fetch them directly from memory.
@@
 Actual compilation times will vary with different CUDA versions and hardware quality, particularly disk performance.
-On modern devices, the compilation overhead can be reduced in various ways.
-In practice, especially in homogeneous cloud environments, it can be advantageous to store compiled kernels on shared or remote storage and download them as needed.
-The similar approach has been used in the machine learning community, e.g.
-PyTorch AOTInductor.
-Alternatively, retaining intermediate .ptx files and loading them directly allows users to bypass one step of compilation.
-Integral pattern Algorithm
-Compilation
-time (ms)
-Binary size
+In practice, the compilation overhead can be reduced in various ways.
+Especially in homogeneous cloud environments, it can be advantageous to store compiled kernels on shared or remote storage and download them as needed.
+A similar approach is used in the machine learning community, where PyTorch’s AOTInductor pre-compiles GPU kernels and caches the resulting binaries for reuse across runs.
+Alternatively, retaining the intermediate PTX assembly (a CUDA specific compilation intermediate) and loading it directly allows users to skip the final machine-code generation step on subsequent invocations.
+Integral pattern Algorithm Compilation time (ms) Binary size (KB)
 $(ss|ss)/(11|11)$ 1q1t 284.8 25
@@
 The number of significant shell quartets in each JK calculation is identical in both cases.
-In double-precision, JoltQC reproduces GPU4PySCF’s results exactly, so we omit a detailed FP64 accuracy analysis.
+In double precision, JoltQC reproduces GPU4PySCF’s J and K matrices to machine precision for all test systems, confirming that the JIT-compiled kernels are numerically identical to the AOT implementation.
+We therefore focus the accuracy analysis on the single-precision results.
 Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
 We use a customized TeraChem v1.9 build with f-orbital support.
-Although both GPU4PySCF and TeraChem employ the Schwarz inequality to bound integral magnitudes, they apply it at different levels: TeraChem at the primitive shell level and GPU4PySCF at the contracted level.
-Their SCF initial guesses and DIIS settings also differ, so our GPU4PySCF–TeraChem comparisons are not strictly apples-to-apples.
-For all calculations shown, the screening threshold for both J and K matrices is set to $10^{-13}$.
+Before presenting timings, we note several differences that affect the fairness of comparisons.
+GPU4PySCF and JoltQC use identical screening (Schwarz inequality at the contracted level), so their JK kernel timings are directly comparable.
+TeraChem applies Schwarz screening at the primitive shell level, uses different SCF initial guesses and DIIS settings, and converges in fewer iterations (e.g.
+11 vs.
+14 for the gly30 molecule).
+Consequently, GPU4PySCF/JoltQC–TeraChem comparisons reflect differences in both algorithms and settings, and are not strictly apples-to-apples.
+For all calculations, the screening threshold is set to $10^{-13}$.
 SCF with GPU4PySCF/JoltQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations.
@@
 When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both JoltQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations.
+For the cc-pVQZ basis, which involves higher angular momentum, the FP32 energy deviations remain sub-milliHartree (0.16 mHa for tamoxifen and 0.23 mHa for sphingomyelin), indicating that single-precision accuracy does not degrade with increasing basis set size.
 A more practical mixed-precision algorithm is among our future directions.
-Molecule Basis
-GPU4PySCF/FP64
-- JoltQC/FP32
-GPU4PySCF/FP64
-- TeraChem/FP64
-JoltQC/FP32
-- TeraChem/FP32
+Molecule Basis GPU4PySCF/FP64 - JoltQC/FP32 GPU4PySCF/FP64 - TeraChem/FP64 JoltQC/FP32 - TeraChem/FP32
 gly30 6-31G* 0.23 mHa 0.0005 mHa -0.07 mHa
@@
 valinomycin def2-TZVPP 0.68 mHa -0.001 mHa 0.63 mHa
-Table 4: The accuracy of Hartree-Fock (HF) energy with JoltQC v0.1 against TeraChem v1.9. GPU4PySCF/FP64 and JoltQC/FP64 matches exactly and is not shown in the table.
+Tamoxifen $c$-pVQZ 0.16 mHa - -
+Sphingomyelin $c$-pVQZ 0.23 mHa - -
+Table 4: The accuracy of Hartree-Fock (HF) energy with JoltQC v0.1 against TeraChem v1.9.
+GPU4PySCF/FP64 and JoltQC/FP64 matches exactly and is not shown in the table.
+Dashes indicate that the software does not support the required angular momentum (g functions in $c$-pVQZ).
+Each column reports the total energy difference $E_{A}-E_{B}$ between the two codes listed.
 As shown in Table 5, JoltQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis.
 Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal.
-In double-precision, JoltQC outperforms TeraChem by about 4x when def2-TZVPP basis set is used, and a less significant 1.6x for 6-31G* basis.
+In double-precision, JoltQC outperforms TeraChem by about 4x when the def2-TZVPP basis set is used, and a less significant 1.6x for the 6-31G* basis.
+The advantage grows further with larger basis sets.
+For the cc-pVQZ basis, which includes g-type functions not supported by TeraChem v1.9, JoltQC achieves a 7–8× speedup over GPU4PySCF v1.4 on the tamoxifen, raffinose, and sphingomyelin systems — nearly doubling the gain observed with def2-TZVPP.
+This trend confirms that JIT compilation becomes increasingly beneficial as the angular momentum of the basis set rises, precisely the regime where register pressure and loop complexity penalize statically compiled kernels the most.
 In single-precision, JoltQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP.
-Both TeraChem and JoltQC still employ double precision for eigenvalue decomposition and DIIS, making these linear-algebra routines a significant cost on the A10-24 GB GPU.
-Isolating the JK kernel, JoltQC is approximately three times faster than TeraChem.
+Both TeraChem and JoltQC still employ double precision for eigenvalue decomposition and DIIS.
+On the A10-24G, which has only 1/32 the FP64 throughput of its FP32 units, these FP64 linear-algebra routines consume a disproportionate fraction of the total SCF time, diminishing the effective speedup from FP32 integral evaluation.
+Isolating the JK kernel on the A10-24G with FP32 arithmetic, JoltQC is approximately three times faster than TeraChem across the benchmark systems (Table 5, JK/cycle column).
 Molecule Protocol GPU4PySCF TeraChem JoltQC
 JK/cycle SCF JK/cycle SCF JK/cycle SCF
-gly30 HF/6-31G*/A100/FP64 1.9 s 28.1 s 1.3 s 19.1 s 0.8 s 10.4 s
-gly30 HF/def2-TZVPP/A100/FP64 78.4 s 1,125.7 s 91.4 s 1,055.1 s 18.8 s 268.1 s
-gly30 HF/6-31G*/A10/FP32 - - 1.0 s 23.1 s 0.7 s 16.5 s
-gly30 HF/def2-TZVPP/A10/FP32 - - 63.5 s 870.8 s 20.0 s 406.1 s
+gly30 6-31G*/A100/FP64 1.9 s 28.1 s 1.3 s 19.1 s 0.8 s 10.4 s
+gly30 def2-TZVPP/A100/FP64 78.4 s 1,125.7 s 91.4 s 1,055.1 s 18.8 s 268.1 s
+gly30 6-31G*/A10/FP32 - - 1.0 s 23.1 s 0.7 s 16.5 s
+gly30 def2-TZVPP/A10/FP32 - - 63.5 s 870.8 s 20.0 s 406.1 s
+valinomycin 6-31G*/A100/FP64 2.4 s 34.5 s 1.9 s 24.2 s 1.1 s 14.8 s
+valinomycin def2-TZVPP/A100/FP64 113.0 s 1,579.6 s 162.0 s 1,641.6 s 30.0 s 419.2 s
+valinomycin $c$-pVTZ/A100/FP64 96.0 s 1,349.5 s 140.0 s 1,534.5 s 24.0 s 330.6 s
+tamoxifen $c$-pVQZ/A100/FP64 152.0 s 1,977.4 s - - 21.4 s 278.9 s
+raffinose $c$-pVQZ/A100/FP64 188.0 s 2,252.9 s - - 28.6 s 343.6 s
+sphingomyelin $c$-pVQZ/A100/FP64 227.0 s 2,726.0 s - - 29.7 s 356.1 s
 Table 5: Benchmark among GPU4PySCF v1.4, TeraChem v1.9, and JoltQC v0.1, on A100-80G using FP64 arithmetic, and on A10-24G using FP32 arithmetic.
+Dashes indicate that the software does not support the required angular momentum (g functions in $c$-pVQZ)
 5 Conclusion
-We optimized the low–angular-momentum electron repulsion integral algorithm in GPU4PySCF v1.4 using just-in-time (JIT) compilation and introduced a novel high–angular-momentum integral method.
-The core CUDA implementation for both algorithms spans only 1,000 lines of code in total.
-JIT compilation not only accelerates the evaluation of Gaussian-type-orbital integrals but also provides developers with greater flexibility.
-Our streamlined implementation delivers a 2-4x performance improvement over GPU4PySCF v1.4 and is significantly easier to maintain.
-Additionally, it supports single- and double-precision arithmetic, accepts multiple density matrices as input, and integrates seamlessly with the existing GPU4PySCF framework.
+We have demonstrated that just-in-time compilation is a powerful paradigm for accelerating electron repulsion integral evaluation on GPUs.
+By making angular momentum, orbital contraction pattern, and primitive counts compile-time constants, the JIT compiler can aggressively unroll loops, eliminate branches, and optimize register allocation — yielding a 2× speedup for small basis sets (6-31G*) and up to 4× for large basis sets (def2-TZVPP) over GPU4PySCF v1.4 on an A100-80G GPU.
+In single precision on an A10-24G, JoltQC is approximately three times faster than TeraChem v1.9 for the JK kernels.
+Beyond performance, JIT compilation fundamentally changes the development model: the core NVRTC implementation spans only 1,000 lines for both algorithms, compared to $\sim$20,000 lines in GPU4PySCF v1.4.
+This compactness is not incidental — it is a direct consequence of delegating specialization to the compiler rather than encoding it by hand.
+The resulting code is easier to maintain, extend, and adapt to new hardware.
+We expect JIT compilation to become a standard tool in the quantum-chemistry software stack as GPU architectures continue to diversify.
+We release JoltQC as an open-source library to serve as both a practical tool and a foundation for applying JIT compilation more broadly in quantum chemistry.
+JoltQC supports single- and double-precision arithmetic, accepts multiple density matrices, and integrates seamlessly with GPU4PySCF.
 6 Future Work and Limitations
@@
 A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels.
-The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post Hartree-Fock methods, and a wide array of molecular properties are all essential.
+The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post-Hartree-Fock methods, and a wide array of molecular properties are all essential.
 By integrating JoltQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
@@
 Other research efforts [18, 21] show that 90% of the computations can be in FP32.
-A naive mixed precision strategy can be easily proposed based on this work.
-One should notice, even though the value of integral can be represented in FP32, not every underlying arithmetic operation for computing the integral can necessarily be computed with FP32.
-More effort is needed to rigorously implement a strategy without error accumulation.
+A straightforward mixed-precision strategy follows directly from the Schwarz inequality: shell quartets whose integral upper bound exceeds the FP32 representable range are computed in FP64, while the remaining 90% are computed in FP32.
+JoltQC’s JIT framework makes this trivial to implement — the precision is already a template parameter (Table I), so both FP32 and FP64 kernels are compiled and dispatched based on the screening estimate.
+A key subtlety is that even when the final integral value lies within FP32 representable range, the intermediate arithmetic — particularly the recursive evaluation of Rys polynomials and the accumulation of primitive contributions — may involve catastrophic cancellation or large intermediate values that require FP64.
+A rigorous mixed-precision strategy must therefore identify which computational steps can safely use FP32 and which must remain in FP64 to avoid error accumulation.
+This is a nontrivial compiler-level analysis that we leave to future work.
 Screening techniques.
@@
 For example, the MBIE [15], QQR [19], and CSAM [28].
-MBIE provides a rigorous upperbound, but it is expensive to calcuate.
-QQR and CSAM take advantage of distance dependence in their estimates, hence improving the performance for large systems.
-But those estimates are less rigorous.
+MBIE provides a rigorous upperbound, but it is expensive to calculate.
+QQR and CSAM exploit distance dependence in their estimates, offering tighter bounds for large systems at the cost of reduced rigor compared to the Schwarz inequality.
+Integrating these screening strategies into JoltQC is straightforward: because the screening is performed on the host before kernel dispatch, it is independent of the JIT-compiled kernels and can be swapped without recompilation.
 Author Contributions